OpenStudy (anonymous):

solve by graphing 3x+4y=20 3x-2y=8 can someone help me

OpenStudy (anonymous):

To graph any line, you should first find out x-intercept and y-intercept. x-intercept is where value of y is zero in the equation. y-intercept is where value of x is zero in the equation. Let us take 3x+4y=20 x-intercept is 20/3 (where y=0) y-intercept is 20/4=5 (where x=0)

OpenStudy (anonymous):